OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 4x − 7 = 29 for x.

OpenStudy (anonymous):

4x=29+7=36 x=9

OpenStudy (lilg132):

4x - 7 = 29 4x = 29 + 7 4x = 36 4x / 4 = 36 / 4 x = 9 remember to click good answer

OpenStudy (anonymous):

that correct and a good answer